### v1

# ONEBench to Test Them All: Sample-Level Benchmarking Over Open-Ended Capabilities

Abstract
Traditional fixed test datasets fall short in evaluating the open-ended capabilities of foundation models. To address this limitation, we propose ONEBench (OpeN-Ended Benchmarking), a new testing paradigm that consolidates individual evaluation datasets into a unified, ever-expanding sample pool. ONEBench allows users to generate custom, open-ended evaluation benchmarks from this sample pool, corresponding to specific capabilities of interest. By aggregating and reusing samples across test sets, ONEBench enables the assessment of diverse capabilities beyond those covered by the original test sets, while mitigating overfitting and dataset bias. Most importantly, it frames model evaluation as a collective process of selecting and aggregating sample-level tests.
[…]
1 Introduction
ONEBench to rule them all (evals),
ONEBench to find them (samples),
ONEBench to test them all (models),
and by queried ability rank them.
Deep learning has arrived in the post-dataset era [^1] From a talk by Alexei Efros at ICML 2020. With the rapidly expanding range of zero-shot capabilities of foundation models, the focus of model evaluation has moved beyond singular, dataset-specific performance measurements obtained by splitting a fixed collection of data into training and test sets. Instead, foundation models are employed as general knowledge and reasoning engines across all domains in which they prove to be useful. This creates a pressing need to characterize their open-ended set of capabilities across various metrics in zero-shot settings [39], for flexible evaluation over several use-cases. However, traditional static benchmarks, which test generalization on fixed test splits, are unable to effectively probe the ever-evolving set of capabilities of foundation models. This raises an important question: How can benchmarking adapt to measure an open-ended set of capabilities?
We propose a solution based on dynamic sample-level evaluation, which we call ONEBench for OpeN-Ended Benchmarking, where test sets for particular capabilities are generated ad-hoc from a large pool of individual annotated data samples. These sample-level evaluations act as atomic units of measurement that can be combined into an exponential variety of aggregations. Due to this flexibility, the sample pool and corresponding annotation metrics can be continuously updated to include new evaluations. Additionally, this approach can reduce dataset bias—systematic quirks in the data arising from how it was collected [151, 87, 171]. Finally, by combining samples across test sets, ONEBench can better capture real-world diversity [111].
The most important feature of ONEBench is the potential to democratize evaluation. Unlike traditional benchmarks, typically created by individual groups based on their own criteria for data collection and evaluation procedures [27, 7], ONEBench allow the integration of test sets from multiple sources reflecting a wide range of perspectives, use-cases, and objectives. This flexibility allows different interest groups with varying needs to collaboratively define their own evaluations by selecting the most appropriate combination of tests to suit their specific requirements. Moreover, the design of ONEBench challenges the dominant approach of chasing single benchmark scores, which do not indicate the difficulty of individual data instances [32], in favour of a plurality of rankings and dynamic, granular, multi-faceted evaluation.
Challenges in ONEBench. To build ONEBench, we must address two main challenges: (a) Heterogeneity and (b) Incompleteness. Heterogeneity means that model measurements span different metric types, such as binary (correct/incorrect), numeric (BLEU scores), and ordinal (preference rankings). This diversity makes it difficult to aggregate measurements and standardize comparisons across different models. Incompleteness, on the other hand, arises from models being evaluated on different, unequal subsets of testing data, rendering direct aggregation unfair and inaccurate. Traditional benchmarks use a multi-task setting, where each component test set requires evaluating all models over a fixed set of samples using a single metric, completely sidestepping both these issues.
Figure 1: The ONEBench Framework. Left: ONEBench comprises a set of models, a pool of data samples spanning multiple test sets, metadata describing models and data samples, and a collection of heterogeneous, sample-level measurements. Right: the user formulates a query to capture the desired model capability, using a mix of structured metadata filters and semantic search. Selected models are then ranked on a subset of data samples that meet the specified criteria.
Solution and Theoretical Guarantees. To tackle these challenges, we apply social choice theory, treating data samples as voters who express preferences among models. By converting all measurements into ordinal rankings, we leverage well-established principles to create a robust algorithm for aggregating over heterogeneous and incomplete data. Our approach assumes a random utility model based on the Plackett-Luce framework, which provides guarantees for accurately recovering ground-truth utility scores from input measurements. This approach ensures that our model rankings are both theoretically sound and practical, with rapid convergence guarantees enabling accurate rankings from relatively small amounts of data.
Empirical Validation. We develop two instantiations of ONEBench: ONEBench-LLM for language models and ONEBench-LMM for multimodal(vision and language) models. These benchmarks unify evaluations across their respective domains by aggregating data from diverse sources, from arena-style human preference data [18, 91] to heterogeneous multi-task leaderboards [8, 79, 176, 23]. Our empirical results demonstrate that the Plackett-Luce model [119, 93] is a good fit for aggregating real-world benchmarks, showing a high correlation with ground-truth score-based rankings over homogeneous datasets. Importantly, we demonstrate that this strong correlation holds even when up to $95$% of the data is missing. This robustness allows us to reduce costs by 2020{20}20 times with little loss in performance by selecting a subset of data samples for evaluation. Finally, we compare Plackett-Luce rankings with widely adopted ranking metrics like ELO [30] and Bradley-Terry [15] and show that our method outperforms them on accuracy and robustness to missing information.
Personalized Aggregation. Consider the following scenario: you are a scientist in a biochemistry lab and require an LLM to assist with designing experiments related to antibodies. With ONEBench, you can input a query, such as “immunology” or “antibodies”, and receive a dynamically constructed benchmark that ranks models based on their performance on this specific capability. While the optimal selection of personalized capability sets is an emerging research field, we present a proof of concept by distinguishing between tasks (e.g., reading comprehension) and concepts (e.g., Clostridium bacteria). We show how a combination of structured filters and flexible semantic search allows users to define their capability of interest with respect to these two dimensions and perform targeted evaluations resulting in personalized model rankings.
In essence, ONEBench represents a democratized, open-source collection of diverse evaluation samples enriched with detailed metadata, driven by an aggregation method designed to robustly rank models across heterogeneous metrics and incomplete evaluation data. Users can conduct semantic searches and apply structured query filters to dynamically generate a benchmark tailored to their specific use case. They can also contribute new evaluation samples and sample-level model measurements, which can then be instantly aggregated, producing personalized rankings. This framework enables lifelong aggregation of arbitrary test sets with unprecedented flexibility and precision.
2 ONEBench: Formulation
At the heart of ONEBench is the idea of homogenizing performance evaluation across benchmarks by replacing benchmark-specific metrics with rankings. Importantly, this can be done at the level of individual data samples. In the following, we describe the process of construction and evaluation in detail, together with mathematical guarantees.
2.1 Components
The goal of building ONEBench (${\mathcal{B}}$) from a growing set of benchmarks $({\mathcal{B}}_{k})_{k=1}^{N}$ is to evaluate a collection of models (${\mathcal{M}}$) using an ever-expanding test pool of data instances ($\mathcal{D}$) which may be annotated with additional meta-data specifying the capabilities ($\mathcal{C}$) tested. To cope with the diversity of data originating from different benchmarks, sample-level rankings (${\mathcal{S}}$) are created for all data instances in the test pool. We provide a schematic overview of ONEBench in Fig. 1 and describe each component below:
i) Pool of Data. ${\mathcal{D}}{=}{((x_{1},y_{1}),\dots,(x_{n},y_{n}))}$ denotes an ordered collection of test data instances $x_{i}$ with annotation $y_{i}$. An example of a data instance $x_{k}$ is the question “What was the dominant strain of Flu in 2010? Select among the four choices.” with the reference answer “H1N1/09” represented by $y_{k}$. In addition, information about tested capabilities can be provided as metadata, for example as a list of keywords such as temporal Q&A, pandemics, history, biology, virology, multiple-choice Q&A, beyond the specific dataset it originates from. Typically, the data samples are obtained via pooling from $N$ different benchmarks $({\mathcal{B}}_{k})_{k=1}^{N}$ and we refer to the subset of data instances obtained from benchmark ${\mathcal{B}}_{k}$ as ${\mathcal{D}}_{k}\subseteq\mathcal{D}$.
ii) Models. ${\mathcal{M}}{=}{(f_{base},f_{1},\dots,f_{m})}$ is a set of $m+1$ models, whose capabilities are evaluated with respect to a baseline $f_{base}$. An example of $f_{base}$ is a random model. The original benchmarks $({\mathcal{B}}_{k})_{k=1}^{N}$ will likely cover different sets of models ${\mathcal{M}_{\mathcal{B}_{k}}}\subseteq\mathcal{M}$ for their evaluations.
iii) Sample-level Rankings. For each data instance $(x_{j},y_{j})\in\mathcal{D}$ a sample-level ranking ${s_{j}\in\mathcal{S}}$ is created for the subsets of models ${\mathcal{M}_{j}:=\mathcal{M}\cap\mathcal{M}_{\mathcal{B}_{k(j)}}}$ where $k(j)$ denotes the index of the benchmark from which the data instance $(x_{j},y_{j})$ was collected. Importantly, sample-level rankings are a function of the metrics used by the different benchmarks that discard any information about the specifics of the metrics. This is the key to our approach of enabling the aggregation across heterogeneous evaluation paradigms and metrics. More specifically, $s_{j}\in\mathcal{S}$ represents an ordinal ranking over the models $\mathcal{M}_{j}$ for sample $(x_{j},y_{j})$ represented by a permutation $\sigma_{j}$ such that $f_{\sigma_{j}(1)}\succeq\dots\succeq f_{\sigma_{j}(m_{j})}$ where $m_{j}=|\mathcal{M}_{j}|$ is the number of models compared in the $j$-th sample-level ranking. In addition, for each $k$ we distinguish the case $f_{\sigma(k-1)}\succ f_{\sigma(k)}$ if $f_{\sigma(k-1)}$ performs better than $f_{\sigma(k)}$ and $f_{\sigma(k-1)}\sim f_{\sigma(k)}$ in case of indistinguishable performance. Thus, each sample-level ranking $s_{j}\in\mathcal{S}$ can be uniquely determined by a mapping $\sigma_{j}:\{1,\dots,m_{j}\}\to\{1,\dots,m\}$ with $\sigma_{j}(k)$ providing the index of the model in $\mathcal{M}$ that is on the $k$-th place in the ordering for the $j$-th sample-level ranking and $\pi_{j}\in\{\succ,\sim\}^{m_{j}-1}$ defining the corresponding binary sequence of pairwise performance relations.
Ordinal Rankings and Information Loss. Using ordinal measurements leads to information loss, which can impede downstream aggregation algorithms due to the data processing inequality ([148], Section 2.8). This principle asserts that any estimation made from processed data cannot outperform estimation based on the original, unprocessed data. However, cardinal measurements frequently suffer from calibration issues, even within a single metric [134]. Consequently, in practice, ordinal measurements can paradoxically outperform cardinal ones despite the inherent information loss.
iv) Capabilities. To support the selective retrieval of all relevant sample-level rankings in $\mathcal{B}$ based on user queries, it is possible to endow the sample-level rankings with additional capability $c\in\mathcal{C}$. Of course, modelling the range of capabilities that different evaluators might be interested in is a research challenge in itself. Here, we only provide a proof of concept, for which we define two categories of capabilities: (1) tasks, like multiple-choice question answering, captioning, and translation, and (2) concepts like makeup, geometry, tarantula, etc. The rationale for this broad interpretation in ONEBench is to explore which capabilities can be reliably evaluated in a dynamic, flexible manner through queries. Since the set of capabilities is open-ended, we only tag the data samples with task information. The relevant samples for a given concept are retrieved from ONEBench dynamically at test time through semantic search.
Lifelong Expansion of ONEBench. The data pool ($\mathcal{D}$) and model IDs ($\mathcal{M}$) are stored as a table, while sample-level model measurement ($\mathcal{S}$) are represented as a relational database linking these two tables. Constructing a lifelong heterogenous benchmark augments $\mathcal{D}$, $\mathcal{M}$ and $\mathcal{S}$ with the following operations: ${\mathcal{B}}{=}{(\mathcal{D},\mathcal{M},\mathcal{S},\text{{insert}}_{\mathcal{D}},\text{{insert}}_{\mathcal{M}}},\text{{insert}}_{\mathcal{S}}$). Operations $\text{{insert}}_{\mathcal{D}}$ and $\text{{insert}}_{\mathcal{M}}$ for expanding the data pool are straightforward: add new samples and new models to the corresponding table. The $\text{{insert}}_{\mathcal{S}}$ operation inserts a sample-level ranking. Additional metadata is saved to enable retrieval over database rows sharing the same original metric, such as ‘BLEU score’, or ‘exact match’.
2.2 Querying Capabilities for Personalized Evaluation
To evaluate a given capability, ONEBench takes a dynamic approach. First, we retrieve ($\texttt{retrieve}_{\mathcal{D}}$) samples that match the query. Then, we aggregate ($\texttt{aggregate}_{\mathcal{S},\mathcal{D}}$) the sample-level rankings to produce the overall ranking.
Retrieve ($\texttt{retrieve}_{\mathcal{D}}$). In this step, the system selects relevant data instances based on a user’s query. The query language is flexible and allows retrieving data instances that semantically relate to a specific topic or match certain criteria. The retrieval is implemented through a combination of k-nearest neighbors (kNN) search on dense embeddings using the query as the input and structured queries that take advantage of the unified data schema. We provide extensive empirical analysis to test retrieval quality.
Aggregate ($\texttt{Aggregate}_{\mathcal{S},\mathcal{D}}$). The sample-level measurements over the retrieved subset of data samples is combined using the random utility modelling approach [163], which defines a joint probability distribution over all measurements assuming statistical independence:
$p(s_{1},\dots,s_{n_{\infty}}|\gamma_{1},\dots,\gamma_{m})=\prod_{j=1}^{n_{\infty}}p(s_{j}=[.]_{(\sigma_{j},\pi_{j})}|\gamma_{1},\dots,\gamma_{m}).$
The Placket-Luce framework assumes the following probability model:
$p\left(s_{j}=[.]_{(\sigma_{j},\pi_{j})}\right)=\frac{\gamma_{\sigma_{j}(1)}}{\underbrace{\sum_{k=1}^{m_{j}}\gamma_{\sigma_{j}(k)}}_{f_{\sigma_{j}(1)}}}\times\frac{\gamma_{\sigma_{j}(2)}}{\underbrace{\sum_{k=2}^{m_{j}}\gamma_{\sigma_{j}(k)}}_{f_{\sigma_{j}(2)}}}\times\cdots\times\frac{\gamma_{\sigma_{j}(m_{j}-1)}}{\underbrace{\gamma_{\sigma_{j}(m_{j}-1)}+\gamma_{\sigma_{j}(m_{j})}}_{f_{\sigma_{j}(m_{j})}}},$
defining one parameter $\gamma_{k}$ for each model $f_{k}$ that determines its performance relative to all other models. To aggregate the model performances over all sample-level rankings, we determine the parameters
$\hat{\gamma}_{1},\dots\hat{\gamma}_{m}=\operatorname*{argmax}_{(\gamma_{1},\dots\gamma_{m})\in\mathbb{R}^{m}}\log p(s_{1},\dots,s_{n_{\infty}}|\gamma_{1},\dots,\gamma_{m})$
with maximum likelihood estimation. The global ranking is given by the permutation $\sigma_{\infty}$ for which $\hat{\gamma}_{\sigma_{\infty}(1)}>\dots>\hat{\gamma}_{\sigma_{\infty}(m)}$. The maximum likelihood condition uniquely determines all performance parameters $\{\hat{\gamma}_{k}\}_{k=1}^{m}$, as the likelihood function is strictly concave. The parameters of the Plackett-Luce model are identifiable up to an arbitrary additive constant. Consistency and asymptotic normality can also be shown under certain assumptions about the comparison graph [50]. We refer to the estimated latent variables $\{\hat{\gamma}_{k}\}_{k=1}^{m}$ as model scores. A model with a higher score is more likely to perform better on a randomly picked sample-level task than one with a lower score. To fix the arbitrary additive constant, we set the score of the baseline model $\hat{\gamma}_{baseline}$ to zero.
3 Aggregation in ONEBench: Theory and Practice
We view aggregating sparse ordinal preferences over models through a computational social choice lens—samples are voters, models are candidates, and the aggregation algorithm is the voting mechanisms [16]. Using established methods, we aggregate ordinal comparisons with partial data to produce a global ranking and analyze the properties of this resultant ranking.
3.1 Theoretical Foundations
We begin by postulating a ground-truth statistical model generating the data, which is converted into ordinal comparisons ($\mathcal{S}$) [^2] This contrasts with Zhang and Hardt [174], who view aggregation as classical voting, analysing tradeoffs in aggregating voter preferences rather than uncover an underlying ranking.. Specifically, we use a random-utility model [149], where each model $f_{i}$ is associated with a utility distribution $\mathcal{U}_{f_{i}}$. Preferences between models $f_{i}$ and $f_{j}$ are based on comparing sampled utilities, i.e., $f_{i}\prec f_{j}:=u(f_{i})<u(f_{j})$, where $u_{f}\sim\mathcal{U}_{f}$. Since computing maximum likelihood estimates over general random-utility models is computationally hard [163], we focus on the Plackett–Luce model [119, 94], the only known exception that allows for tractable maximum likelihood estimates (MLE).
Property 1: Identifiability. We first ask: Are the utility distributions for all models recoverable? The Plackett-Luce model allows identifying the utility distribution (up to an arbitrary additive constant) if all models are compared via a directed path [57, 163] [^3] Recall that using the reference model $f_{\text{base}}$ removes the additive ambiguity.. Consistency and asymptotic normality hold under specific assumptions about the comparison graph [50].
Property 2: Sample-Efficient Convergence from Sparse Data. Given that identifiability is asymptotic, we then ask: How sample-efficient is the algorithm for recovering the utility distribution? With partial rankings of size $k$, the MLE is surprisingly sample efficient while being minmax-optimal [46, 102]. Specifically, sampling $k$ model comparisons from the model set $|\mathcal{F}|$ independently and uniformly at random for $|\mathcal{D}|$ samples induces an expander graph with high probability, which provides guarantees on sample-efficiency of recovery, with $|\mathcal{D}|=\nicefrac{{\Omega(|\mathcal{F}|)}}{{k}}$ samples being necessary, and $|\mathcal{D}|=\nicefrac{{\Omega(|\mathcal{F}|\log|\mathcal{F}|)}}{{k}}$ samples being sufficient. Efficient algorithms like those in Agarwal et al. [1] and Maystre and Grossglauser [102] achieve these bounds. Rank-breaking techniques, used in our empirical evaluation, also offer near-optimal solutions [141].
Property 3: Active Aggregation. In ONEBench, we can strategically select model comparisons, by framing selection as an online multi-armed bandit problem. Sample efficiency can be improved with PAC guarantees [145, 124, 126], significantly outperforming random comparisons [103].
Property 4: Social Properties. The Plackett-Luce model ensures computational efficiency and recoverability of the underlying ranking. However, to design democratic systems for decision-making, it is essential also to have fair aggregation. Ensuring fairness involves trade-offs[174], because different notions of fairness often conflict, and, depending on the intended application areas, differing or even opposing preferences may be valid [6, 10, 38]. The Plackett-Luce model offers “procedural fairness” [84] (Section 2.2), satisfying the following three criteria:
  • Anonymity. All voters (samples) are treated equally, ensuring the system does not rely on a single vote. The rankings remain unchanged if the input sample set is permuted.
  • Neutrality. The ranking is invariant to the identities of the models, ensuring fairness among alternatives. This means permuting the models similarly permutes the new ranking.
  • Independence from Irrelevant Alternatives (IIA). The relative ranking of two models is unaffected by other alternatives in a given sample, as guaranteed by Luce’s axiom of choice [93]. This provides grounding for incomplete model evaluations.
3.2 Translating Theory to Practice
Figure 2: Top-10 model ranking changes across different aggregation methods. Our Plackett-Luce-based rank aggregation method (Ours) shows the most similarity to the Ground Truth model rankings (GT). However, there is a progressive degradation in ranking accuracy for LMArena (LMArena) and Elo (ELO). Comparisons are shown for ONEBench-LLM (top) and ONEBench-LMM (bottom). Our method best preserves the ranking of the top-10 models.
We now empirically validate our framework, aiming to show that: (i) the Plackett-Luce model works well on real-world data, (ii) our aggregation method is sample-efficient, and (iii) it handles high levels of incompleteness. Below, we describe our setup and address these points.
3.2.1 Setup
Benchmarks. We conduct experiments using four popular benchmarks with established model rankings based on average scores specific to the benchmark: HELM [79] and Open LLM Leaderboard [8] for LLMs, and VHELM [23] and LMMs-Eval [176] for LMMs. We use the score-based rankings over individual datasets as ground truth and define our sample pool as the sum of the data pools of all consituent datasets of a given benchmark. To test the faithfulness of our aggregation strategy we compare the resulting rankings to the original leaderboards. These leaderboards evaluate foundation models across varied tasks with different metrics, serving as good indicators of real-world performance.
Ground Truth. The current system of benchmarking involves evaluating models on individual test sets and measuring the mean score per model. This holds even for benchmarks that combine test sets such as HELM [79], VHELM [23], etc. We consider these scores as the ground truth measurement and generate a ground truth model ranking from these scores. Since we aggregate benchmarks spanning multiple measurement metrics, we implement a min-max normalization of numeric measurements to bring all benchmark samples to the same 0-1 score range. Our final ground truth refers to the model rankings derived from the mean score across all benchmarks.
Methods. We evaluate three model ranking methods:
(i) Elo Score [30]: A competitive game rating system adapted to rank models through pairwise comparisons, adjusting scores based on wins or losses to reflect win-rate reliability.
(ii) LMArena Ranking [18]: A method for LLM ranking using the Bradley-Terry model [15], which estimates model rankings through Maximum Likelihood Estimation (MLE) based on pairwise comparisons using an underlying ELO model.
(iii) Our Method: Our approach leverages the Plackett-Luce model [102] to aggregate pairwise comparisons using partial rank breaking [141], which achieves a significant speed-up for rank estimation methods.
Metrics. We compare the rankings generated by each method to the ground-truth from the leaderboards using Kendall’s $\tau$, a standard correlation metric for rankings. Each method is tested three times, and we report the mean and variance. We additionally check that the top-$k$ models are reliably recovered.
3.2.2 Is Plackett-Luce a Good Fit for Real-World Data?
Metric HELM Open LLM VHELM LMMs-Eval
Leaderboard
Elo Score 0.347 ± 0.132 0.213 ± 0.065 0.639 ± 0.024 0.332 ± 0.109
LMArena Ranking 0.852 ± 0.001 0.969 ± 0.000 0.697 ± 0.000 0.427 ± 0.000
Our Method 0.877 ± 0.001 0.997 ± 0.000 0.799 ± 0.000 0.641 ± 0.000
Table 1: Kendall’s $\tau$ correlations to ground-truth ranking for different aggregation algorithms. Results show improvements over ELO and LMArena rankings, with notable correlation boosts on ONEBench-LMM leaderboards, including LMMs-Eval (41.65%) and VHELM (14.63%).
Figure 3: Sample-efficient convergence (top) and robustness to sparsity. Kendall $\tau$ between ground-truth ranking and different ranking methods as random individual data samples are dropped (top) and model measurements are randomly removed (bottom). Methods typically remain robust to missing data, with Plackett-Luce consistently achieving higher correlation, even with 95% measurements missing.
Q1. Is it a good fit? We assess whether the Plackett-Luce model performs well on large-scale benchmark data by comparing our aggregation algorithm’s rankings to the leaderboard rankings. As shown in Table 1, our algorithm achieves a high positive Kendall’s $\tau$, indicating strong alignment with the ground truth rankings.
Q2. Is it better than current metrics? In addition to evaluating fit, we also compare our method to popular algorithms like Elo and LMArena. Table 1 shows that our algorithm consistently outperforms these methods, demonstrating its superior performance for large real-world datasets.
Q3. Are the top-k models preserved? For practitioners, the critical concern is whether the top models are ranked correctly. Figure 2 shows that our algorithm effectively preserves the top-10 model rankings compared to ground truth, while outperforming state-of-the-art methods in maintaining accurate top-k rankings.
Conclusion. The Plackett-Luce model fits real-world data well, outperforming other methods in both overall Kendall’s $\tau$ and top-10 model rankings, proving its effectiveness for large-scale benchmarks. The underlying reason is that we avoid the limitations of Elo-based methods, which rely on assumptions that do not apply to foundation models [13].
3.2.3 Sample Efficiency and Handling Incomplete Rankings
We now empirically test the sample efficiency and robustness to incomplete data of our framework.
Q1. Is Our Algorithm Sample-Efficient? We systematically reduce the number of samples and re-rank the models using various methods, calculating Kendall’s $\tau$ for each. Missing data is simulated from 0% to 99%, with 10% intervals until 90%, followed by 1% increments. As shown in Fig. 3, our method maintains stable performance even with up to 95% samples missing, demonstrating that it can achieve accurate rankings with far fewer data points—up to 20x less than current benchmarks.
Q2. Can our Algorithm Aggregate Highly Sparse Rankings? We evaluate the method’s ability to handle highly incomplete data by removing model comparisons for each sample and re-ranking the models. We randomly remove a given fraction of model measurements from each sample and re-rank using the three aggregation methods. Again, we simulate data removal from 0% to 99%, with increments as before. As shown in Fig. 3, our method performs well even with 95% fewer model comparisons, proving it can recover accurate rankings with highly sparse data. This is crucial for ONEBench, where models cannot be expected to be evaluated on the entire data pool.
Conclusion. Our algorithm provides significant sample efficiency, maintaining accurate rankings with 20x fewer data points, and is robust to highly sparse input rankings.
4 ONEBench: Creation & Capability Querying
Figure 4: Constituent datasets of ONEBench-LLM (left) and OneBench-LMM (right). We provide task type, metric, and license about each dataset in Tab. 2 and Tab. 3.
After evaluating the robustness of our aggregation method across incomplete and heterogeneous measurements, we present the overall system applied to two large-scale, real-world instances of ONEBench for foundation models: LLMs and LMMs. We first outline how these benchmarks were created, then show how to use them to test arbitrary capabilities, and finally highlight key insights gained.
Dataset Source Task Size Metric License
Cardinal
LegalBench [43] HELM Legal 1K QEM Unknown
MATH [53] HELM Maths 1K QEM MIT
MedQA [59] HELM Medical 1K QEM MIT
NarrativeQA [68] HELM Openbook QA 1K F1 Apache-2.0
NaturalQuestions [71] HELM Search Engine Queries 1K F1 CC BY-SA 3.0
OpenbookQA [106] HELM Openbook QA 1K EM Apache-2.0
WMT 2014 [12] HELM Machine translation 1K BLEU CC-BY-SA-4.0
ARC [19] Leaderboard General QA $1.1$K EM CC-BY-SA-4.0
HellaSwag [170] Leaderboard Reasoning $10$K EM MIT
TruthfulQA [82] Leaderboard General QA $817$ EM Apache-2.0
Winogrande [129] Leaderboard Reasoning $1.2$K EM Apache-2.0
GSM8K [20] HELM + Leaderboard Maths $1.3$K QEM MIT
MMLU [52] HELM + Leaderboard General QA $13.8$K EM MIT
Ordinal
Chatbot Arena [18] Chatbot Arena Pairwise Battles $51$K - CC BY 4.0
Table 2: Datasets in ONEBench-LLM. A diverse collection of benchmarks testing the abilities of LLMs in areas such as law, medicine, mathematics, question answering, reasoning and instruction following, as well as the performance of LLMs in pairwise battles.
Dataset Source Task Size Metric License
Cardinal
A-OKVQA [133] VHELM VQA $7.2$K QEM Apache-2.0
Bingo [24] VHELM Bias+Hallucination $886$ ROUGE Unknown
Crossmodal-3600 [147] VHELM Captioning $1.5$K ROUGE CC BY-SA 4.0
Hateful Memes [64] VHELM Hate Speech $1$K QEM Custom(Meta)
Mementos [160] VHELM Sequential Reasoning $945$ GPT CC-BY-SA-4.0
MultipanelVQA [33] VHELM VQA $200$ QEM MIT
OODCV-VQA [152] VHELM VQA $1$K QEM CC-BY-NC-4.0
PAIRS [35] VHELM Bias $508$ QEM Unknown
Sketchy-VQA [152] VHELM VQA $1$K QEM CC-BY-NC-4.0
AI2D [61] LMMs-Eval Maths+Science $3.09$K QEM Apache-2.0
IconQA [88] LMMs-Eval Docs and Infographics $43$K ANLS CC BY-SA 4.0
InfoVQA [101] LMMs-Eval Docs and Infographics $6.1$K ANLS Unknown
LLaVA-in-the-Wild [85] LMMs-Eval Multi-disciplinary $60$ GPT4 Apache-2.0
ChartQA [99] LMMs-Eval Docs and Infographics $2.5$K QEM GPL-3.0
CMMMU [173] LMMs-Eval Multi-disciplinary $900$ QEM CC-BY-4.0
DocVQA [100] LMMs-Eval Docs and Infographics $10.5$K ANLS Unknown
MMBench [86] LMMs-Eval Multi-disciplinary $24$K GPT Apache-2.0
MMVET [167] LMMs-Eval Multi-disciplinary $218$ GPT Apache-2.0
MP-DocVQA [150] LMMs-Eval Docs and Infographics $5.2$K QEM MIT
NoCaps [3] LMMs-Eval Captioning $4.5$K ROUGE MIT
OK-VQA [98] LMMs-Eval VQA $5.1$K ANLS Unknown
RefCOCO [60, 97] LMMs-Eval Captioning $38$K ROUGE Apache-2.0
ScienceQA [89] LMMs-Eval Maths+Science $12.6$K EM CC BY-NC-SA 4.0
TextCaps [139] LMMs-Eval Captioning $3.2$K ROUGE CC BY 4.0
TextVQA [140] LMMs-Eval VQA $5$K EM CC BY 4.0
COCO [83] VHELM+LMMs-Eval Captioning $45.5$K ROUGE CC-BY-4.0
Flickr30k [165] VHELM+LMMs-Eval Captioning $31$K ROUGE CC-0 Public Domain
GQA[56] VHELM+LMMs-Eval Scene Understanding $12.6$K QEM CC-BY-4.0
MathVista [90] VHELM+LMMs-Eval Maths+Science $1$K QEM/GPT4 CC-BY-SA-4.0
MME [36] VHELM+LMMs-Eval Multi-disciplinary $2.4$K QEM/C+P Unknown
MMMU [169] VHELM+LMMs-Eval Multi-disciplinary $900$ QEM CC BY-SA 4.0
POPE [78] VHELM+LMMs-Eval Hallucination $9$K QEM/EM MIT
SEED-Bench [73, 74] VHELM+LMMs-Eval Multi-disciplinary $42.5$K QEM/EM Apache
VizWiz [44] VHELM+LMMs-Eval VQA $4.3$K QEM/EM CC BY 4.0
VQAv2 [42] VHELM+LMMs-Eval VQA $214$K QEM/EM CC BY 4.0
Ordinal
Vision Arena [91] - Pairwise Battles $9$K - MIT
LMMs-Eval(Prometheus2) [66] - Pairwise Battles $610$K - MIT
Table 3: Datasets in ONEBench-LMM: a diverse collection of benchmarks testing the abilities of LLMs in tasks such as general VQA, image captioning, hate speech detection, bias and hallucination understanding, maths and science, documents and infographics, scene understanding and sequential reasoning as well as the performance of LMMs in pairwise battles.
4.1 Creation of ONEBench-LLM & ONEBench-LMM
We develop two instantiations of ONEBench: (1) ONEBench-LLM for language models and (2) ONEBench-LMM for multimodal (vision and language) models, as outlined in  Fig. 4.
4.1.1 ONEBench-LLM
Data Pool $\mathcal{D}$. For ONEBench-LLM (Tab. 2), we source data from the Open LLM Leaderboard [8], HELM [79], and LMArena [18]. Open LLM Leaderboard and HELM aggregate several individual benchmarks (such as MMLU [52] and HellaSwag [170]), while LMArena uses pairwise model comparisons based on user-generated prompts, with user votes determining the superior model. Metrics which are converted to sample-wise ordinal rankings here include F1-Score, Exact Match (EM), and Quasi-Exact Match (QEM), as well as pairwise preferences from LMArena.
Models $\mathcal{F}$. For ONEBench-LLM, we use the 100 most downloaded models from Open LLM Leaderboard and all 79 models from HELM (as of v1.9.0), including both proprietary models like GPT-4o [112] and open-weights ones like LLaMA-3 [105]. A full list of evaluated models is provided in Appx. A.
4.1.2 ONEBench-LMM
Data Pool $\mathcal{D}$. For ONEBench-LMM (Tab. 3), data is sourced from VHELM, LMMs-Eval, and WildVisionArena. Similar to ONEBench-LLM, VHELM and LMMs-Eval aggregate individual datasets like MMMU [169] and VQAv2 [42], while WildVisionArena uses pairwise tests for LMMs through image-based chats. We convert a diverse set of metrics to sample-wise rankings, from binary metrics like EM, QEM, to real-valued scores like ROUGE [81], Perception (P), and Cognition (C) scores from MME [36]. We additionally combine pairwise comparisons from WildVisionArena with LLM-As-A-Judge preferences generated using Prometheus-2 [66], which correlate highly with human judgments. The preference comparisons are sampled randomly from LMMs-Eval while avoiding overlap with cardinal measurements.
Models $\mathcal{F}$. For ONEBench-LMM, we use 14 models from LMMs-Eval [176] and 25 models from VHELM [23], including proprietary models like Gemini Pro Vision [146] and open-weights models like LLaVA [85]. A complete list of evaluated models is provided in Appx. A.
4.2 Capabilities and Concept Probing: Results & Insights
Figure 5: Capability Probing (Qualitative): we provide six sample retrieval results for a set of queries covering a diverse set of topics and report the top-5 models for each query.
Benchmark ##\##Concepts Cohen-$\kappa$ mAP CMC@1 CMC@10
ONEBench-LLM 40 0.793 0.8462 0.95 1.0
ONEBench-LMM 50 0.912 0.7337 0.94 0.96
Table 4: Capability Probing (Quantitative): we provide a summary of the number of concepts curated for capability probing, along with the inter-annotator agreement and retrieval metrics.
Here, we present empirical results on generating arbitrary test sets and rankings. Our goal is to enable users to make targeted queries within ONEBench and interactively search across the data pool, helping them identify the best models for their needs. To achieve this, we extend our system with a flexible mechanism for personalized aggregation, allowing users to (1) retrieve relevant data instances through structured filters and semantic search, and (2) dynamically generate model rankings based on the retrieved samples.
Setup. Given a query, the system retrieves relevant data samples using a combination of flexible semantic search and rigid constraints. This concept querying mechanism provides a personalized comparison of foundation model capabilities. We use two querying mechanisms. (i) Semantic search: we perform k-NN lookup in the embedding space of all-MiniLM-L6-v2 [123] for language tasks and SigLIP-B16 [172] for vision-language tasks, using cosine similarity. We retrieve the top-k samples for a given concept with a tuned cut-off similarity score of $0.3$ and $0.7$ for ONEBench-LLM and ONEBench-LMM, respectively. (ii) Metadata search: we verify that per-sample metadata satisfies the constraints defined in the query. Some benchmarks, such as MMMU, are equipped with detailed metadata, including categories like image type (‘diagram’, ‘table’, etc.), question type (‘multiple-choice’), field (‘Electronics’), and subfield (‘Analog Electronics’). Other benchmarks, such as image captioning datasets (e.g. COCO, NoCaps, Flickr30k), contain little metadata.
With these resources, we gather representative samples for individual queries across diverse test sets, and aggregate the ordinal sample-level model rankings using the Plackett-Luce model to produce a model ranking for that particular query.
Concepts Tested. We curated a diverse set of 50 concepts to test the breadth and versatility of ONEBench, ranging from domain-specific knowledge, such as the Coriolis Effect, to broader academic disciplines like Neuroscience, and everyday consumer goods like the Apple iPad. We showcase 6 of them in the main paper and present the rest in Appx. B. We present the results from concept querying in Figure 5 and summarize our insights below:
Insight 1. Are the retrieved data samples accurate? Two expert annotators manually reviewed and filtered out incorrect matches  [^4] The inter-annotator agreement, measured by Cohen’s Kappa, is shown in Table 4, with high values of 0.793 and 0.912 indicating strong consistency between annotators.. To evaluate the quality of the retrieved samples, we report average precision (AP) scores for a random subset of queried concepts in Fig. 5, with a full list of scores in Appx. B. Aggregating over all tested concepts in Table 4, our mAP over the concepts is 0.84 and 0.73 for ONEBench-LLM and ONEBench-LMM, respectively, demonstrating that we can reliably retrieve samples that match the intended capabilities, although there is substantial scope for improvement in some cases. Please refer to the per-concept AP in Tab. 5 for a better indicator of underrepresented concepts in ONEBench. Note that the retrieval mechanism is expected to only improve with better retrieval models, larger test sets that cover a more diverse range of capabilities, and the integration of more sophisticated querying mechanisms in ONEBench.
Insight 2. Do models perform differently across queries? A key check is to verify whether models perform distinctly across different capability queries. If the results are similar regardless of the query, fine-grained querying may be less useful, as the top model from a generic leaderboard could be a good candidate for any specific capability, as is common practice currently. However, we observe in Fig. 5 and Fig. 6 that different models perform well on different domains and concepts. This suggests that ONEBench returns valid candidate models for arbitrary user queries.
5 Related Works
Multi-task Benchmarks as Broad Capability Evaluators. Multi-task leaderboards have been the standard for benchmarking foundation models. Examples include GLUE [158], decaNLP [104], SuperGLUE [157], BigBench [142], Dynabench [65], Open LLM Leaderboard [8], CLIP-Benchmark [72], ELEVATOR [75], StableEval [153] and DataComp-38 [37], as well as massive multitask benchmarks like XTREME [138] and ExT5 [5]. However, concerns have arisen regarding the limitations of multi-task benchmarks [14]. Issues include saturation and subsequent discarding of samples [80, 11, 113, 31, 162], susceptibility to dataset selection [25], obscuring progress by evaluation metrics [22, 132], training on test tasks [154, 28, 109, 107, 143, 159], and data contamination [29, 96, 41, 26, 127, 40, 128]. ONEBench tackles these challenges by enabling extensive reuse of samples for broader model comparisons, avoiding task selection bias through democratized sourcing of samples, and using ordinal rankings to avoid evaluation minutia. Sample-level evaluation with sparse inputs also allows selective removal of contaminated data for fairer comparisons. Moreover, by supporting over-ended, evolving evaluation, it makes it harder to train on all test tasks, as opposed to fixed leaderboards that are easier to game.
On Aggregation across Benchmarks. The dominant approach to benchmarking has traditionally been multi-task benchmarks, where the most common aggregation strategy is the arithmetic mean of scores across individual tasks. However, this approach assumes that the scoring metrics are homogeneous and scaled correctly, and treat tasks of different complexities equally [108, 117]. In consequence, simple normalization preprocessing influences the rankings [21], and makes them nearly entirely dependent on outlier tasks [2]. Simply changing the aggregation method from arithmetic to geometric or harmonic mean can change the ranking [136]. Similarly, including irrelevant alternative models can change statistical significance or even change the ranking entirely [9, 174]. Mean-aggregation also has significant failure modes in handling missing scores in benchmarks [54]. The benchmarking paradigm is hence shifting towards adopting evaluation principles from other fields, such as non-parametric statistics and social choice theory [16, 125]. We use ordinal rankings instead of scores, similar to LMArena. However, unlike Arena, we use the pairwise variant of the Plackett-Luce model, which has been shown to have advantages both theoretically and empirically [116]. We benefit from some of its theoretical properties like identifiability, sample-efficient convergence, provable robustness to irrelevant alternatives, non-dominance of outliers and empirical robustness across a wide range of real-world factors which affect ranking. Moreover, we do not aggregate over benchmarks in the first place—our primary proposal is to avoid monolithic benchmarks and consider aggregation on a sample level, needing to tackle incomplete and heterogeneous measurements. We note that several other social-choice theory-based models such as score-based models [125, 137] based on the Condorcet-winner criterion [164] have been proposed, yet they were primarily applied for aggregation on multi-task benchmarks, whereas a crucial component of our proposal is to break down the benchmark boundaries and aggregate heterogeneous samples.
Dynamic Evaluation and Active Testing. Some previous works like [58, 69, 70, 130, 55, 178] tackle the ‘active testing’ problem, where the goal is to identify small “high-quality” test data-subsets, from a large pool of uncurated evaluation data. These works typically assume that the cost of unlabeled test data acquisition is low whereas the cost of acquiring per-instance labels is high. However, as pointed out in [122], these assumptions are unrealistic for foundation models, as both the acquisition of test data and label annotations can be tedious in general. Hence, in our work, we tackle a broader problem: given a large testing data pool, how can we curate and query to produce a consistent and targeted set of model rankings?
Efficient Evaluation. As evaluation suites have grown, associated inference costs have also increased. Recent research has focused on creating compressed subsets of traditional benchmarks to address this issue [155, 120, 156, 177, 115, 67, 114]. Popular approaches include subsampling benchmarks to preserve correlations with an external source like LMArena [111], sample clustering to gauge sample difficulty and then sub-sampling[156], item-response-theory based methods for informatively sampling a subset of samples for evaluation[120], or designing evolving sample-level benchmarks [122]. While the work of Prabhu et al. [122] is similar to us in principle, it requires binary metrics as input and does not handle incomplete input matrices, which is necessary for aggregation over multiple time steps. We precisely address these limitations by showing efficient evaluation while accommodating incomplete data and extending it to ordinal ranks.
Democratizing Evaluation. Standard image classification and retrieval benchmarks are collected from platforms like Flickr, which are predominantly Western-centric [4, 135]. This has raised the important question: “Progress for whom?”, with many seminal works showcasing large disparities in model performance on concepts [51, 110], tasks [47, 48, 49], and even input samples [45, 121, 144] from the Global South. In response, works have developed benchmarks tailored to diverse cultures and demographics to include their voice in measuring progress [92, 110, 121, 118]. Further works have tried to create personalized, task-specific benchmarks for flexibly evaluating models based on user-preferences [17, 77, 131, 168]— Zhang et al. [175] created Task-Me-Anything that enables users to input specific queries that then get processed to provide model rankings or responses to the query. However, their system is entirely procedurally generated, thereby not reflecting the real-world use-cases that models are typically subjected to in practice. Further, they are restricted to the fixed set of instances in their task generator pool. We take a different approach by creating flexible benchmarks where individuals, and contributing entities, can add their own samples and preferences collected from both real-world benchmarks and live model arenas like LM-Arena, thereby providing users with a realistic overview of model rankings on practical scenarios. Further, during capability testing, users can select similar preferences, making ONEBench more inclusive than traditional test sets.
6 Conclusions and Open Problems
This work tackles scalable benchmarking of arbitrary capabilities of foundation models, requiring a shift from traditional fixed training and test splits, by introducing ONEBench, an open-ended benchmarking framework for foundation models. Our open-source, democratized benchmarking methodology allows diverse evaluation samples and model measurements with detailed metadata. This affords creating customized benchmarks and testing arbitrary capabilities, including using semantic and structured searches. We provide a principled aggregation mechanism, that is both theoretically grounded and empirically validated to be robust to incomplete data and heterogeneous measurements across evaluations. We demonstrate the utility of ONEBench in two domains: ONEBench-LLM and ONEBench-LMM, showing how dynamic probing reveals new insights into model performance on specific tasks, domains, or concepts. This combination of theoretical rigour, empirical results, and practical flexibility makes ONEBench a valuable tool for comprehensively evaluating foundation models. We provide some promising directions for improvement below:
  1. 1.
    Testing Limits and Scaling Up ONEBench: currently, our prototype comprises less than 100K samples in ONEBench-LLM and under 1M in ONEBench-LMM. These pools can be greatly expanded and diversified by expanding to incorporating all existing LLM and LMM benchmarks. Our retrieval mechanisms are designed to scale efficiently as the test pool grows in size and diversity.
  2. 2.
    Exploring Other Aggregation Algorithms: while we use the Plackett-Luce model for aggregating diverse measurements, there exist other algorithms from computational social choice theory with different trade-offs. A comprehensive evaluation of these alternatives could offer new insight for aggregating model performance.
  3. 3.
    Structured Querying and Enhanced Retrieval: One can improve retrieval by better querying mechanisms using models like ColBERT [62] and ColPALI [34], further optimized using DSPy [63]. A particularly interesting direction is allowing compositional queries, where users combine multiple queries to test behaviour in foundation models, similar to works like ConceptMix [161] and SkillMix [166].
  4. 4.
    On the Limits of Capability Probing: While we currently allow broad, open-ended inputs to probe capabilities, some are easier to assess than others [95, 76]. As foundation models become more generalizable, a thorough analysis identifying which capabilities can be easily, reliably evaluated, which are possible to evaluate but challenging, and which are in principle impossible to evaluate is needed—this will help improve benchmarking effectiveness.
Acknowledgements
The authors would like to thank (in alphabetic order): Nikhil Chandak, Karel D’Oosterlinck, Shashwat Goel, Shyamgopal Karthik, Bálint Mucsányi and Tim Weiland for their helpful suggestions and feedback. Special thanks to Bo Li for (promptly and consistently) providing sample-level model evaluation logs of LMMs-Eval, Yujie Lu for providing early access to the battle data of Vision Arena and Palzer Lama for creating the logo of ONEBench and helping with figures. The authors would also like to thank Federico D’Agostino for recommending the title of the paper and the name of the benchmark. AP and MB acknowledge financial support via the Open Philanthropy Foundation funded by the Good Ventures Foundation. VU and SD thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) and the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. VU is supported by a Google PhD Fellowship in Machine Intelligence. MB acknowledges support via the CRC 1233 on Robust Vision and from the Machine Learning Cluster of Excellence, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645.
Appendix A Models used in ONEBench:Further Details
In this section, we provide a deeper insight into the models used in the creation of ONEBench. It is important to note that ONEBench-LLM and ONEBench-LMM have complementary characteristics: while ONEBench-LLM has fewer data samples $\mathcal{D}_{k}$, they are evaluated on more models $\mathcal{M}_{k}$, while ONEBench-LMM contains (significantly) more data samples but they are evaluated on less models.
A.1 ONEBench-LLM: Open LLM Leaderboard
The Open LLM Leaderboard [8] was created to track progress of LLMs in the open-source community by evaluating models on the same data samples and setup for more reproducible results and a trustworthy leaderboard where all open-sourced LLMs could be ranked.
However, due to the abundance of models found on the leaderboard and the lack of adequate documentation, and therefore reliability, of many of these models being evaluated, we rank the models based on the number of downloads, as a metric of adoption of these models by the community. We provide the total list of models as an artefact and list the top 100 models below:
  1. 1.
    01-ai/Yi-34B-200K
  2. 2.
    AI-Sweden-Models/gpt-sw3-126m
  3. 3.
    BioMistral/BioMistral-7B
  4. 4.
    CohereForAI/c4ai-command-r-plus
  5. 5.
    CohereForAI/c4ai-command-r-v01
  6. 6.
    Deci/DeciLM-7B-instruct
  7. 7.
    EleutherAI/llemma__\__7b
  8. 8.
    EleutherAI/pythia-410m
  9. 9.
    Felladrin/Llama-160M-Chat-v1
  10. 10.
    Felladrin/Llama-68M-Chat-v1
  11. 11.
    FreedomIntelligence/AceGPT-7B
  12. 12.
    GritLM/GritLM-7B
  13. 13.
    Intel/neural-chat-7b-v3-1
  14. 14.
    JackFram/llama-160m
  15. 15.
    Nexusflow/NexusRaven-V2-13B
  16. 16.
    Nexusflow/Starling-LM-7B-beta
  17. 17.
    NousResearch/Hermes-2-Pro-Mistral-7B
  18. 18.
    NousResearch/Meta-Llama-3-8B-Instruct
  19. 19.
    NousResearch/Nous-Hermes-2-Mixtral-8x7B-DPO
  20. 20.
    NousResearch/Nous-Hermes-2-SOLAR-10.7B
  21. 21.
    NousResearch/Nous-Hermes-2-Yi-34B
  22. 22.
    OpenPipe/mistral-ft-optimized-1227
  23. 23.
    Qwen/Qwen1.5-0.5B
  24. 24.
    Qwen/Qwen1.5-0.5B-Chat
  25. 25.
    Qwen/Qwen1.5-1.8B
  26. 26.
    Qwen/Qwen1.5-1.8B-Chat
  27. 27.
    Qwen/Qwen1.5-110B-Chat
  28. 28.
    Qwen/Qwen1.5-14B
  29. 29.
    Qwen/Qwen1.5-14B-Chat
  30. 30.
    Qwen/Qwen1.5-32B-Chat
  31. 31.
    Qwen/Qwen1.5-4B
  32. 32.
    Qwen/Qwen1.5-4B-Chat
  33. 33.
    Qwen/Qwen1.5-72B-Chat
  34. 34.
    Qwen/Qwen1.5-7B
  35. 35.
    Qwen/Qwen1.5-7B-Chat
  36. 36.
    SeaLLMs/SeaLLM-7B-v2
  37. 37.
    TinyLlama/TinyLlama-1.1B-Chat-v1.0
  38. 38.
    TinyLlama/TinyLlama-1.1B-intermediate-step-1431k-3T
  39. 39.
    VAGOsolutions/SauerkrautLM-Mixtral-8x7B-Instruct
  40. 40.
    abhishekchohan/mistral-7B-forest-dpo
  41. 41.
    ahxt/LiteLlama-460M-1T
  42. 42.
    ai-forever/mGPT
  43. 43.
    alignment-handbook/zephyr-7b-sft-full
  44. 44.
    augmxnt/shisa-gamma-7b-v1
  45. 45.
    bigcode/starcoder2-15b
  46. 46.
    bigcode/starcoder2-3b
  47. 47.
    bigcode/starcoder2-7b
  48. 48.
    cloudyu/Mixtral__\__7Bx4__\__MOE__\__24B
  49. 49.
    codellama/CodeLlama-70b-Instruct-hf
  50. 50.
    cognitivecomputations/dolphin-2.2.1-mistral-7b
  51. 51.
    cognitivecomputations/dolphin-2.6-mistral-7b-dpo
  52. 52.
    cognitivecomputations/dolphin-2.9-llama3-8b
  53. 53.
    daekeun-ml/phi-2-ko-v0.1
  54. 54.
    deepseek-ai/deepseek-coder-1.3b-instruct
  55. 55.
    deepseek-ai/deepseek-coder-6.7b-base
  56. 56.
    deepseek-ai/deepseek-coder-6.7b-instruct
  57. 57.
    deepseek-ai/deepseek-coder-7b-instruct-v1.5
  58. 58.
    deepseek-ai/deepseek-math-7b-base
  59. 59.
    deepseek-ai/deepseek-math-7b-instruct
  60. 60.
    deepseek-ai/deepseek-math-7b-rl
  61. 61.
    google/codegemma-7b-it
  62. 62.
    google/gemma-1.1-7b-it
  63. 63.
    google/gemma-2b
  64. 64.
    google/gemma-2b-it
  65. 65.
    google/gemma-7b
  66. 66.
    google/gemma-7b-it
  67. 67.
    google/recurrentgemma-2b-it
  68. 68.
    h2oai/h2o-danube2-1.8b-chat
  69. 69.
    hfl/chinese-alpaca-2-13b
  70. 70.
    ibm/merlinite-7b
  71. 71.
    meta-llama/Meta-Llama-3-70B
  72. 72.
    meta-llama/Meta-Llama-3-70B-Instruct
  73. 73.
    meta-llama/Meta-Llama-3-8B
  74. 74.
    meta-llama/Meta-Llama-3-8B-Instruct
  75. 75.
    meta-math/MetaMath-Mistral-7B
  76. 76.
    microsoft/Orca-2-7b
  77. 77.
    microsoft/phi-2
  78. 78.
    mistral-community/Mistral-7B-v0.2
  79. 79.
    mistral-community/Mixtral-8x22B-v0.1
  80. 80.
    mistralai/Mistral-7B-Instruct-v0.2
  81. 81.
    mistralai/Mixtral-8x22B-Instruct-v0.1
  82. 82.
    mistralai/Mixtral-8x7B-Instruct-v0.1
  83. 83.
    mistralai/Mixtral-8x7B-v0.1
  84. 84.
    openai-community/gpt2
  85. 85.
    openai-community/gpt2-large
  86. 86.
    openchat/openchat-3.5-0106
  87. 87.
    openchat/openchat-3.5-1210
  88. 88.
    openchat/openchat__\__3.5
  89. 89.
    sarvamai/OpenHathi-7B-Hi-v0.1-Base
  90. 90.
    speakleash/Bielik-7B-Instruct-v0.1
  91. 91.
    speakleash/Bielik-7B-v0.1
  92. 92.
    stabilityai/stablelm-2-1__\__6b
  93. 93.
    stabilityai/stablelm-2-zephyr-1__\__6b
  94. 94.
    stabilityai/stablelm-zephyr-3b
  95. 95.
    teknium/OpenHermes-2.5-Mistral-7B
  96. 96.
    tokyotech-llm/Swallow-70b-instruct-hf
  97. 97.
    upstage/SOLAR-10.7B-Instruct-v1.0
  98. 98.
    upstage/SOLAR-10.7B-v1.0
  99. 99.
    wenbopan/Faro-Yi-9B
  100. 100.
    yanolja/EEVE-Korean-Instruct-10.8B-v1.0
A.2 ONEBench-LLM: HELM
Similar to the Open LLM Leaderboard, the goal of HELM was to provide a uniform evaluation of language models over a vast set of data samples (termed as scenarios in Liang et al. [79]). HELM, however, has a broader scope of models used for evaluation, employing open, limited-access, and closed models. All models currently used in ONEBench-LLM is listed below:
  1. 1.
    01-ai__\__yi-34b
  2. 2.
    01-ai__\__yi-6b
  3. 3.
    01-ai__\__yi-large-preview
  4. 4.
    ai21__\__j2-grande
  5. 5.
    ai21__\__j2-jumbo
  6. 6.
    ai21__\__jamba-1.5-large
  7. 7.
    ai21__\__jamba-1.5-mini
  8. 8.
    ai21__\__jamba-instruct
  9. 9.
    AlephAlpha__\__luminous-base
  10. 10.
    AlephAlpha__\__luminous-extended
  11. 11.
    AlephAlpha__\__luminous-supreme
  12. 12.
    allenai__\__olmo-7b
  13. 13.
    anthropic__\__claude-2.0
  14. 14.
    anthropic__\__claude-2.1
  15. 15.
    anthropic__\__claude-3-5-sonnet-20240620
  16. 16.
    anthropic__\__claude-3-haiku-20240307
  17. 17.
    anthropic__\__claude-3-opus-20240229
  18. 18.
    anthropic__\__claude-3-sonnet-20240229
  19. 19.
    anthropic__\__claude-instant-1.2
  20. 20.
    anthropic__\__claude-instant-v1
  21. 21.
    anthropic__\__claude-v1.3
  22. 22.
    cohere__\__command
  23. 23.
    cohere__\__command-light
  24. 24.
    cohere__\__command-r
  25. 25.
    cohere__\__command-r-plus
  26. 26.
    databricks__\__dbrx-instruct
  27. 27.
    deepseek-ai__\__deepseek-llm-67b-chat
  28. 28.
    google__\__gemini-1.0-pro-001
  29. 29.
    google__\__gemini-1.0-pro-002
  30. 30.
    google__\__gemini-1.5-flash-001
  31. 31.
    google__\__gemini-1.5-pro-001
  32. 32.
    google__\__gemini-1.5-pro-preview-0409
  33. 33.
    google__\__gemma-2-9b-it
  34. 34.
    google__\__gemma-2-27b-it
  35. 35.
    google__\__gemma-7b
  36. 36.
    google__\__text-bison@001
  37. 37.
    google__\__text-unicorn@001
  38. 38.
    meta__\__llama-2-7b
  39. 39.
    meta__\__llama-2-13b
  40. 40.
    meta__\__llama-2-70b
  41. 41.
    meta__\__llama-3-8b
  42. 42.
    meta__\__llama-3-70b
  43. 43.
    meta__\__llama-3.1-8b-instruct-turbo
  44. 44.
    meta__\__llama-3.1-70b-instruct-turbo
  45. 45.
    meta__\__llama-3.1-405b-instruct-turbo
  46. 46.
    meta__\__llama-65b
  47. 47.
    microsoft__\__phi-2
  48. 48.
    microsoft__\__phi-3-medium-4k-instruct
  49. 49.
    microsoft__\__phi-3-small-8k-instruct
  50. 50.
    mistralai__\__mistral-7b-instruct-v0.3
  51. 51.
    mistralai__\__mistral-7b-v0.1
  52. 52.
    mistralai__\__mistral-large-2402
  53. 53.
    mistralai__\__mistral-large-2407
  54. 54.
    mistralai__\__mistral-medium-2312
  55. 55.
    mistralai__\__mistral-small-2402
  56. 56.
    mistralai__\__mixtral-8x7b-32kseqlen
  57. 57.
    mistralai__\__mixtral-8x22b
  58. 58.
    mistralai__\__open-mistral-nemo-2407
  59. 59.
    nvidia__\__nemotron-4-340b-instruct
  60. 60.
    openai__\__gpt-3.5-turbo-0613
  61. 61.
    openai__\__gpt-4-0613
  62. 62.
    openai__\__gpt-4-1106-preview
  63. 63.
    openai__\__gpt-4-turbo-2024-04-09
  64. 64.
    openai__\__gpt-4o-2024-05-13
  65. 65.
    openai__\__gpt-4o-mini-2024-07-18
  66. 66.
    openai__\__text-davinci-002
  67. 67.
    openai__\__text-davinci-003
  68. 68.
    qwen__\__qwen1.5-7b
  69. 69.
    qwen__\__qwen1.5-14b
  70. 70.
    qwen__\__qwen1.5-32b
  71. 71.
    qwen__\__qwen1.5-72b
  72. 72.
    qwen__\__qwen1.5-110b-chat
  73. 73.
    qwen__\__qwen2-72b-instruct
  74. 74.
    snowflake__\__snowflake-arctic-instruct
  75. 75.
    tiiuae__\__falcon-7b
  76. 76.
    tiiuae__\__falcon-40b
  77. 77.
    writer__\__palmyra-x-004
  78. 78.
    writer__\__palmyra-x-v2
  79. 79.
    writer__\__palmyra-x-v3
A.3 ONEBench-LMM: LMMs-Eval
LMMs-Eval is the first comprehensive large-scale evaluation benchmark for Large Multimodal models, meant “to promote transparent and reproducible evaluations” [176]. The models supported by LMMs-Eval are primarily open-sourced and the full list of currently used models are listed below:
  1. 1.
    idefics2-8b
  2. 2.
    internlm-xcomposer2-4khd-7b
  3. 3.
    instructblip-vicuna-7b
  4. 4.
    instructblip-vicuna-13b
  5. 5.
    internVL-Chat-V1-5
  6. 6.
    llava-13b
  7. 7.
    llava-1.6-13b
  8. 8.
    llava-1.6-34b
  9. 9.
    llava-1.6-mistral-7b
  10. 10.
    llava-1.6-vicuna-13b
  11. 11.
    llava-1.6-vicuna-7b
  12. 12.
    llava-7b
  13. 13.
    llava-next-72b
  14. 14.
    qwen__\__vl__\__chat
A.4 ONEBench-LMM: VHELM
Finally, ONEBench-LMM comprises VHELM, an extension of HELM for Vision-Language models. The models currently used by us, spanning open, limited-access, and closed models, are as follows:
  1. 1.
    anthropic__\__claude__\__3__\__haiku__\__20240307
  2. 2.
    anthropic__\__claude__\__3__\__opus__\__20240229
  3. 3.
    anthropic__\__claude__\__3__\__sonnet__\__20240229
  4. 4.
    google__\__gemini__\__1.0__\__pro__\__vision__\__001
  5. 5.
    google__\__gemini__\__1.5__\__pro__\__preview__\__0409
  6. 6.
    google__\__gemini__\__pro__\__vision
  7. 7.
    google__\__paligemma__\__3b__\__mix__\__448
  8. 8.
    huggingfacem4__\__idefics2__\__8b
  9. 9.
    huggingfacem4__\__idefics__\__80b
  10. 10.
    huggingfacem4__\__idefics__\__80b__\__instruct
  11. 11.
    huggingfacem4__\__idefics__\__9b
  12. 12.
    huggingfacem4__\__idefics__\__9b__\__instruct
  13. 13.
    llava__\__1.6__\__mistral__\__7b
  14. 14.
    llava__\__1.6__\__vicuna__\__13b
  15. 15.
    llava__\__1.6__\__vicuna__\__7b
  16. 16.
    microsoft__\__llava__\__1.5__\__13b__\__hf
  17. 17.
    microsoft__\__llava__\__1.5__\__7b__\__hf
  18. 18.
    mistralai__\__bakllava__\__v1__\__hf
  19. 19.
    openai__\__gpt__\__4__\__1106__\__vision__\__preview
  20. 20.
    openai__\__gpt__\__4__\__vision__\__preview
  21. 21.
    openai__\__gpt__\__4o__\__2024__\__05__\__13
  22. 22.
    openflamingo__\__openflamingo__\__9b__\__vitl__\__mpt7b
  23. 23.
    qwen__\__qwen__\__vl
  24. 24.
    qwen__\__qwen__\__vl__\__chat
  25. 25.
    writer__\__palmyra__\__vision__\__003
Appendix B Capability Testing Across Arbitrary Queries
B.1 Queries: List and Additional Results
Concept ONEBench-LLM AP ONEBench-LMM AP
Common Queries
apple ipad 0.7435 0.1985
architecture 0.7683 0.8981
beach 0.7152 0.5698
biochemistry 0.9778 0.7303
boat 0.7728 0.8829
botany 0.9876 0.7556
bus 0.9035 0.9739
car 0.9140 0.8477
cell(biology) 0.9937 0.5075
china tourism 0.6392 1.0000
cigarette advertisment 0.7249 0.6590
coffee maker 0.8426 0.4057
components of a bridge 0.9222 0.5865
decomposition of benzene(organic chemistry) 0.6745 0.7623
epidemiology 0.9316 0.7991
kirchoffs law(electrical engineering) 0.6572 0.4824
food chain 0.5405 1.0000
game of football 0.8221 1.0000
german shepherd (dog) 0.9359 0.3078
gothic style (architecture) 0.7829 1.0000
law 0.8566 0.4138
literary classics 0.9869 1.0000
macroeconomics 1.0000 0.9570
makeup 1.0000 0.2247
microwave oven 0.7979 1.0000
neuroscience components 0.9844 0.2854
pasta 0.5678 0.2142
perfume 0.5996 0.6355
photosynthesis 0.9848 0.3665
plants 1.0000 0.6488
political diplomacy 0.9529 0.9561
python code 0.8850 0.9444
renaissance painting 0.9270 0.9799
shareholder report 1.0000 0.8317
sheet music 0.8322 0.9750
solar cell battery 0.8853 0.8082
thermodynamics 0.9567 0.8852
united states of america 0.8096 0.8642
vaccines 0.8572 0.3411
volanic eruption 0.7905 0.9229
Queries testing Visual Capabilities
bike leaning against a wall - 0.8271
child playing baseball - 0.9638
coriolis effect - 0.7063
dijkstras shortest path algorithm - 0.9135
empty bridge overlooking the sea - 0.5934
judo wrestling - 0.6092
man in a suit - 0.5611
musical concert - 0.9879
sine wave - 0.4232
woman holding an umbrella - 0.8821
Table 5: Aggregate Average Precision(AP) for ONEBench-LLM and ONEBench-LMM concepts.
Figure 6: Additional qualitative analysis for ONEBench’s capability probing for selected queries.